Given the formula A=2*pi*r(r+h). Solve for h.
Answer by addingup(3677)

A=2*pi*r(r+h). Solve for h
:
2*pi*r(r+h) = A
r+h = A/(2*pi*r)
Now solve for h. You can do it, subtract r from both sides and you're done.
John